# Optimal sampling of tensor networks targeting wave function’s fast decaying tails

### v1

@@ Abstract
In the era of the early development of quantum processors and simulators Monroe (2002); Monroe and Kim (2013); Saffman (2018); Wendin (2017); Bruzewicz et al. (2019); Adams et al. (2019); Kaushal et al. (2020); Kjaergaard et al. (2020); Morgado and Whitlock (2021), a high demand arises for numerical emulators of programmable quantum devices: classical algorithms capable of replicating the input, real-time processing, and ouput of a quantum machine at small and intermediate scales Nagy and Savona (2019); Qiskit contributors (2023); The cuQuantum development team (2023); Efthymiou et al. (2022). Such demand is motivated by benchmarking and certification of the programmable quantum device Jaschke et al. (2022) as well as exploring the extent of quantum advantage: classical simulation of quantum machines has shown remarkable progress in recent experiments Arute et al. (2019); Zhong et al. (2020) and new numerical algorithms have been developed to overcome existing obstacles Zhou et al. (2020); Liu et al. (2021); Bulmer et al. (2022); Daley et al. (2022); Oh et al. (2023).
@@
Tensor networks are often employed to numerically characterize intermediate-scale quantum systems and devices. As they store spatial correlations as compressed information, they are an invaluable tool for describing many-body quantum systems with limited entanglement content. Various classes of tensor networks have been developed so far, each adapted to a specific correlation geometry: from matrix product states (MPS) White (1992); Rommer and Östlund (1997, 1999); Schollwöck (2005); Perez-Garcia et al. (2007); McCulloch (2007); Schollwöck (2011) via tree tensor networks (TTN) Shi et al. (2006); Tagliacozzo et al. (2009); Murg et al. (2010); Gerster et al. (2014) to projected entangled pair states (PEPS) F. et al. (2008); Orús (2014), and more Zwolak and Vidal (2004); Verstraete et al. (2004); Rizzi et al. (2008); Cincio et al. (2008); McCulloch (2008); Jordan et al. (2008); Evenbly and Vidal (2009, 2014); Phien et al. (2015); Werner et al. (2016); Silvi et al. (2019); Montangero (2018); Jaschke et al. (2018); Reinić (2022); Arceci et al. (2022); Bañuls (2023); Okunishi et al. (2022). Also in the context of sampling, tensor network methods have emerged as a powerful tool: Indeed, tensor network allow efficient output data acquisition from quantum states and processes Gomez et al. (2022); Yang et al. (2023); Dymarsky and Pavlenko (2022). Sampling of tensor network states has also been used to determine quantities that are otherwise difficult to measure, such quantum state magic Lami and Collura (2023), entanglement characterization Notarnicola et al. (2023), or to sample from distributions via quantum states Wild et al. (2021). Despite considerable recent advancements Han et al. (2018); Bravyi et al. (2022), we identified a sampling strategy meant to outpower known methods.
@@
In this work, we present an efficient sampling procedure, which we label Optimal tEnsor network Sampling (OPES), applicable to any isometric tensor network state Zaletel and Pollmann (2020). The algorithm itself falls into the category of inverse transform sampling algorithms Niederreiter ; Devroye (2006); Givens and Hoeting (2012); as additional challenge in comparison to the inverse transform sampling, we have no initial knowledge of the cumulative probability distribution or its inverse, which even by the end is known only partially for the many-body quantum state. Our approach exploits the capability of isometric tensor networks to efficiently acquire exact conditional probabilities for partial outcome strings, and carefully avoids re-sampling of previously explored outcomes. Optimal tensor network sampling method is thus particularly suited to characterize peaked, non-flat distributions, where we observe a speedup beyond one order of magnitude over the standard sampling. The speedup to be expected depends foremost on the probability distribution, then on the number of samples or coverage to be achieved. To demonstrate the usefulness of the approach, we consider two popular classes of isometric tensor networks, namely MPS and TTN. We identify a substantial speedup both for sampling a fixed number of states as well as for covering a target portion of the probability space; while the exact probability of non-sampled space is known at all times. The OPES algorithm benefits from storing in memory all the partial tensor network contractions, which characterize the conditional probabilities, evaluated until that instant, to avoid repeating calculations. We analyze the advantage offered by the OPES method in five practical examples from different quantum applications, such as image representation, condensed matter or measurement of quantum magic. The examples include pure and mixed states. The developed code is open source and included in the Quantum TEA Ballarin et al. (2024a) library.
@@
This manuscript is structured as follows. First, we focus on the idea underlying OPES in Sec. I, then on the numerical implementations in Secs. II and III. In Sec. IV, we introduce classes of different quantum states, each targeting different applications; we sample with the two methods from these states and analyze the speedup that one can expect in each application. Finally, we summarize our results in Sec. V. The appendices contain the calculation of the exponential speedup of the scaling in App. .1, the full algorithm in App. .2, as well as additional implementation strategies for tensor networks, see App. .3.
I Theory and standard sampling from tensor network states
@@
Although we compare exclusively the performance of the standard algorithm and the OPES algorithm for tensor networks in the following, the motivation originates in experimental realizations and how sampling works there. Figure 1a illustrates how experiments sample from the underlying probability distribution by running multiple shots for three qubits. The ratio of the number of shots for a state $\alpha$ over the total number of shots converges to the exact distribution in the limit of large number of shots. Due to the central limit theorem Fischer (2011), the histogram approximates the actual probability distribution. The basis states $\alpha$ forming the projection in $z$-direction are an intuitive choice for the experiment, which for example can be sampled in an atom-qubit experiment with a fluorescence snapshot Monroe (2002); Saffman (2018); Morgado and Whitlock (2021). Similarly to the experimental results, both tensor network sampling methods can generate the number of shots per state $\alpha$.
@@
|$\ket{\psi}$ $=\sum_{q_{0},\dots,q_{n-1}=0}^{1}c_{q_{0},\dots,q_{n-1}}\ket{q_{0}\dots q_{n-1}}=\sum_{\alpha=0}^{2^{n}-1}c_{\alpha}\ket{\alpha}\,.$ (1)
@@
The boundaries of each $\delta_{\alpha}$ are well defined within the complete probability interval $(0,1)$. Thus, the left boundary $b^{l}_{\alpha}$ of the probability interval associated with the state $\alpha$ can be defined as reported in Eq. (3). We stress that, as shown in Eq. (2), the right boundary is just $b^{r}_{\alpha}=b^{l}_{\alpha}+|c_{\alpha}|^{2}$. Inverse transform sampling algorithms can be used if these intervals are all known Niederreiter ; Devroye (2006); Givens and Hoeting (2012). Then, one can sample a uniform random number in the interval $(0,1)$, which serves as value on the $y$-axis of Fig. 1c, and finds the corresponding state $\alpha$ on the $x$-axis via the inverse cumulative probability distribution. While the inverse cumulative probability distribution is easily accessible once all $p_{\alpha}$ are known, they are initially unknown in the tensor network scenario and the number of states $\alpha$ grows exponentially in system size for many-body systems. But we can already anticipate based on Eq. (3) that neither all $\{|c_{\beta}|^{2}\}_{\beta\neq\alpha}$ nor even $\{|c_{\beta}|^{2}\}_{\beta<\alpha}$ need to be known. Partial knowledge can be sufficient to attribute a random number to the corresponding interval and its state $\alpha$, e.g., knowing the Eq. (3) for an intermediate state with less than $n$ qubits, i.e., the intervals $\delta_{\alpha^{\prime}}$ with $\alpha^{\prime}$ being a subsystem of $\alpha$ containing the first $m<n$ qubits; we explain the details in Sec. III.
The standard approach to sample outcomes $\left(|c_{\alpha}|^{2},\alpha\right)$ is the generation of bitstrings shot by shot. For each shot, the standard method consecutively projects the qubits into their basis state Nielsen and Chuang (2007). As we calculate probabilities shot-by-shot, we obtain the shot count for each bitstring $\alpha$ at the end of the standard sampling as shown in Fig. 1d. Instead of calculating the probability as ratio of shot count over number of shots, tensor networks have direct access to a pair of probabilities $p_{\alpha}$ with each shot. The pair always differs only in the measurement outcome of the last qubit. Measuring the system qubit by qubit, we accumulate a product of conditional probabilities
@@
The fundamental idea of OPES is highlighted in Fig. 1e and f; details of OPES follow in Sec. III. Figure 1e shows the complete knowledge of all probabilities after four shots. Because we can gain the probability of two states with each measurement of a tensor network, the optimal number of necessary measurements scales linearly with the size of the Hilbert space, i.e., as $\mathcal{O}(2^{n-1})$. To reach this scaling, we permanently remove the sampled outcome strings from the probability measure space in a way that they cannot be sampled again, as shown in Fig. 1f. After each step, we thus remove two intervals. Effectively, we renormalize the sampling probabilities of the remaining, missing outcomes. Thus, the sampling probability of a not yet sampled string increases at each shot. Let us suppose we just sampled the couple $(|c_{\alpha^{*}}|^{2},\alpha^{*})$ with probability $p_{\alpha^{*}}$ as the first sample. Then, the probabilities of sampling any $\alpha\neq\alpha^{*}$ at the following iteration are
@@
We observe an exponential speedup reducing the scaling for sampling from $\mathcal{O}(\epsilon^{-1})$ to $\mathcal{O}(\log(\epsilon^{-1}))$ for a decaying probability distribution, with $\epsilon^{-1}$ being the cumulative unknown probability. For details, see App. .1.
II Standard tensor network sampling
@@
Afterward, we repeat the procedure for the second qubit $q_{1}$ with the random number $u_{1}$ computing the conditional probability $p(q_{1}=\ket{0}|q_{0}=\ket{0})=p_{0|0}$ of measuring $q_{1}$ in the state |$\ket{0}$ conditioned on the measurement on $q_{0}$. In the two-qubit example of Fig. 2a, we have all the information to calculate the probability of the sampled state $p_{\ket{01}}$; moreover, we know the probability $p_{\ket{11}}=p_{1|1}p_{0}$ which depends on the same sequence of projections on the previous qubit, or qubits in the general case. We repeat this procedure until the end of the qubit chain, obtaining the probability of the state $\alpha$ as already pointed out in Eq. (4). For a detailed description of how these quantities are efficiently computed via reduced density matrices in an isometric tensor network, refer to App. .3.
III Optimal tensor network sampling
@@
    Furthermore, we highlight that at each iteration we actually compute two probability intervals, i.e. $\delta_{q_{0}q_{1}\dots q_{n-2}q_{n-1}}$ and $\delta_{q_{0}q_{1}\dots q_{n-2}\overline{q_{n-1}}}$.
@@
    Remove the measured probability interval from our sampling interval, i.e. $\Delta=\Delta\setminus\delta_{\alpha}$. Each new sampled number $u_{j}~{}\sim U(\Delta)$ explores a different probability interval $\delta_{\beta\neq\alpha}$.
This presentation of the algorithm is pedagogical for understanding the procedure and generates one sample. Steps 1) to 5) represent in the complete algorithm only a single superiteration with a number of random numbers sampled $n_{\rm r}=1$; we report the complete algorithm in Appendix .2. The main differences are that for efficient computation of the intervals $\delta_{\alpha}$, multiple random numbers $\{u_{j}\}_{j=1,\dots,n_{r}}$ are drawn for each superiteration, repeating then $n_{\rm\scriptscriptstyle S}$ superiterations. While we do not obtain anymore one new state $\alpha$ for each random number sampled from $U(\Delta)$, we are able to cache intermediate results to optimize the procedure; nonetheless, we are still avoiding additional computation for the same interval. For more information about the caching strategy using the information of the intervals, see Appendix .3.

### v2

@@ Abstract
In the era of the early development of quantum processors and simulators [1, 2, 3, 4, 5, 6, 7, 8, 9], a high demand arises for numerical emulators of programmable quantum devices: classical algorithms capable of replicating the input, real-time processing, and ouput of a quantum machine at small and intermediate scales [10, 11, 12, 13]. Such demand is motivated by benchmarking and certification of the programmable quantum device [14] as well as exploring the extent of quantum advantage: classical simulation of quantum machines has shown remarkable progress in recent experiments [15, 16] and new numerical algorithms have been developed to overcome existing obstacles [17, 18, 19, 20, 21, 22].
@@
Tensor networks are often employed to numerically characterize intermediate-scale quantum systems and devices. As they store spatial correlations as compressed information, they are an invaluable tool for describing many-body quantum systems with limited entanglement content. Various classes of tensor networks have been developed so far, each adapted to a specific correlation geometry: from matrix product states (MPS) [23, 24, 25, 26, 27, 28, 29] via tree tensor networks (TTN) [30, 31, 32, 33] to projected entangled pair states (PEPS) [34, 35], and more [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Also in the context of sampling, tensor network methods have emerged as a powerful tool: Indeed, tensor network allow efficient output data acquisition from quantum states and processes [53, 54, 55]. Sampling of tensor network states has also been used to determine quantities that are otherwise difficult to measure, such quantum state magic [56], entanglement characterization [57], or to sample from distributions via quantum states [58]. Despite considerable recent advancements [59, 60], we identified a sampling strategy meant to outpower known methods.
@@
In this work, we present an efficient sampling procedure, which we label Optimal tEnsor network Sampling (OPES), applicable to any isometric tensor network state [61]. The algorithm itself falls into the category of inverse transform sampling algorithms [62, 63, 64]; as additional challenge in comparison to the inverse transform sampling, we have no initial knowledge of the cumulative probability distribution or its inverse, which even by the end is known only partially for the many-body quantum state. Our approach exploits the capability of isometric tensor networks to efficiently acquire exact conditional probabilities for partial outcome strings, and carefully avoids re-sampling of previously explored outcomes. Optimal tensor network sampling method is thus particularly suited to characterize peaked, non-flat distributions, where we observe a speedup beyond one order of magnitude over the standard sampling. The speedup to be expected depends foremost on the probability distribution, then on the number of samples or coverage to be achieved. To demonstrate the usefulness of the approach, we consider two popular classes of isometric tensor networks, namely MPS and TTN. We identify a substantial speedup both for sampling a fixed number of states as well as for covering a target portion of the probability space; while the exact probability of non-sampled space is known at all times. The OPES algorithm benefits from storing in memory all the partial tensor network contractions, which characterize the conditional probabilities, evaluated until that instant, to avoid repeating calculations. We analyze the advantage offered by the OPES method in five practical examples from different quantum applications, such as image representation, condensed matter or measurement of quantum magic. The examples include pure and mixed states. The developed code is open source and included in the Quantum TEA [65] library.
@@
This manuscript is structured as follows. First, we focus on the idea underlying OPES in Sec. 1, then on the numerical implementations in Secs. 2 and 3. In Sec. 5, we introduce classes of different quantum states, each targeting different applications; we sample with the two methods from these states and analyze the speedup that one can expect in each application. Finally, we summarize our results in Sec. 6. The appendices contain the full algorithm in App. A, as well as additional implementation strategies for tensor networks, see App. B.
1 Theory and standard sampling from tensor network states
@@
Although we compare exclusively the performance of the standard algorithm and the OPES algorithm for tensor networks in the following, the motivation originates in experimental realizations and how sampling works there. Figure 1a illustrates how experiments sample from the underlying probability distribution by running multiple shots for three qubits. The ratio of the number of shots for a state $\alpha$ over the total number of shots converges to the exact distribution in the limit of large number of shots. Due to the central limit theorem [66], the histogram approximates the actual probability distribution. The basis states $\alpha$ forming the projection in $z$-direction are an intuitive choice for the experiment, which for example can be sampled in an atom-qubit experiment with a fluorescence snapshot [1, 3, 9]. Similarly to the experimental results, both tensor network sampling methods can generate the number of shots per state $\alpha$.
@@
|$\ket{\psi}$ $=\sum_{q_{0},\dots,q_{n-1}=0}^{1}c_{q_{0},\dots,q_{n-1}}\ket{q_{0}\dots q_{n-1}}$
$=\sum_{\alpha=0}^{2^{n}-1}c_{\alpha}\ket{\alpha}\,.$ (1)
@@
The boundaries of each $\delta_{\alpha}$ are well defined within the complete probability interval $(0,1)$. Thus, the left boundary $b^{l}_{\alpha}$ of the probability interval associated with the state $\alpha$ can be defined as reported in Eq. (3). We stress that, as shown in Eq. (2), the right boundary is just $b^{r}_{\alpha}=b^{l}_{\alpha}+|c_{\alpha}|^{2}$. Inverse transform sampling algorithms can be used if these intervals are all known [62, 63, 64]. Then, one can sample a uniform random number in the interval $(0,1)$, which serves as value on the $y$-axis of Fig. 1c, and finds the corresponding state $\alpha$ on the $x$-axis via the inverse cumulative probability distribution. While the inverse cumulative probability distribution is easily accessible once all $p_{\alpha}$ are known, they are initially unknown in the tensor network scenario and the number of states $\alpha$ grows exponentially in system size for many-body systems. But we can already anticipate based on Eq. (3) that neither all $\{|c_{\beta}|^{2}\}_{\beta\neq\alpha}$ nor even $\{|c_{\beta}|^{2}\}_{\beta<\alpha}$ need to be known. Partial knowledge can be sufficient to attribute a random number to the corresponding interval and its state $\alpha$, e.g., knowing the Eq. (3) for an intermediate state with less than $n$ qubits, i.e., the intervals $\delta_{\alpha^{\prime}}$ with $\alpha^{\prime}$ being a subsystem of $\alpha$ containing the first $m<n$ qubits; we explain the details in Sec. 3.
The standard approach to sample outcomes $\left(|c_{\alpha}|^{2},\alpha\right)$ is the generation of bitstrings shot by shot. For each shot, the standard method consecutively projects the qubits into their basis state [67]. The method can also be used directly on the level of observables [68]. As we calculate probabilities shot-by-shot, we obtain the shot count for each bitstring $\alpha$ at the end of the standard sampling as shown in Fig. 1d. Instead of calculating the probability as ratio of shot count over number of shots, tensor networks have direct access to a pair of probabilities $p_{\alpha}$ with each shot. The pair always differs only in the measurement outcome of the last qubit. Measuring the system qubit by qubit, we accumulate a product of conditional probabilities
@@
The fundamental idea of OPES is highlighted in Fig. 1e and f; details of OPES follow in Sec. 3. Figure 1e shows the complete knowledge of all probabilities after four shots. Because we can gain the probability of two states with each measurement of a tensor network, the optimal number of necessary measurements scales linearly with the size of the Hilbert space, i.e., as $\mathcal{O}(2^{n-1})$. To reach this scaling, we permanently remove the sampled outcome strings from the probability measure space in a way that they cannot be sampled again, as shown in Fig. 1f. After each step, we thus remove two intervals. Effectively, we renormalize the sampling probabilities of the remaining, missing outcomes. Thus, the sampling probability of a not yet sampled string increases at each shot. Let us suppose we just sampled the couple $(|c_{\alpha^{*}}|^{2},\alpha^{*})$ with probability $p_{\alpha^{*}}$ as the first sample. Then, the probabilities of sampling any $\alpha\neq\alpha^{*}$ at the following iteration are
@@
We observe an exponential speedup reducing the scaling for sampling from $\mathcal{O}(\epsilon^{-1})$ to $\mathcal{O}(\log(\epsilon^{-1}))$ for a decaying probability distribution, with $\epsilon^{-1}$ being the cumulative unknown probability. For details, see Sec. 4.
2 Standard tensor network sampling
@@
Afterward, we repeat the procedure for the second qubit $q_{1}$ with the random number $u_{1}$ computing the conditional probability $p(q_{1}=\ket{0}|q_{0}=\ket{0})=p_{0|0}$ of measuring $q_{1}$ in the state |$\ket{0}$ conditioned on the measurement on $q_{0}$. In the two-qubit example of Fig. 2a, we have all the information to calculate the probability of the sampled state $p_{\ket{01}}$; moreover, we know the probability $p_{\ket{11}}=p_{1|1}p_{0}$ which depends on the same sequence of projections on the previous qubit, or qubits in the general case. We repeat this procedure until the end of the qubit chain, obtaining the probability of the state $\alpha$ as already pointed out in Eq. (4). For a detailed description of how these quantities are efficiently computed via reduced density matrices in an isometric tensor network, refer to App. B.
3 Optimal tensor network sampling
@@
    Furthermore, we highlight that at each iteration we actually compute two probability intervals, i.e., $\delta_{q_{0}q_{1}\dots q_{n-2}q_{n-1}}$ and $\delta_{q_{0}q_{1}\dots q_{n-2}\overline{q_{n-1}}}$.
@@
    Remove the measured probability interval from our sampling interval, i.e., $\Delta=\Delta\setminus\delta_{\alpha}$. Each new sampled number $u_{j}~{}\sim U(\Delta)$ explores a different probability interval $\delta_{\beta\neq\alpha}$.
This presentation of the algorithm is pedagogical for understanding the procedure and generates one sample. Steps 1) to 5) represent in the complete algorithm only a single superiteration with a number of random numbers sampled $n_{\rm r}=1$; we report the complete algorithm in Appendix A. The main differences are that for efficient computation of the intervals $\delta_{\alpha}$, multiple random numbers $\{u_{j}\}_{j=1,\dots,n_{r}}$ are drawn for each superiteration, repeating then $n_{\rm\scriptscriptstyle S}$ superiterations. While we do not obtain anymore one new state $\alpha$ for each random number sampled from $U(\Delta)$, we are able to cache intermediate results to optimize the procedure; nonetheless, we are still avoiding additional computation for the same interval. For more information about the caching strategy using the information of the intervals, see Appendix B.